A small electrical motor inside a toy truck moves 4 wheels each 0.10 kg at a speed of 1.1 m/s for 1.5 s. What is the power put out by the motor?


1
Expert's answer
2021-07-13T11:37:56-0400

By the definition of the power we have:


"P=Fv=mav=m\dfrac{v}{t}v=\dfrac{mv^2}{t},""P=\dfrac{4\cdot0.1\ kg\cdot(1.1\ \dfrac{m}{s})^2}{1.5\ s}=0.32\ W."